OpenStudy (anonymous):

Factor the following.. 1. 4r²-4r+1 2. 9s²+12s+4 3. 16t-9 4. 49-100a² 5. 25b²-60b+36 6. 81c²+198c+121

OpenStudy (anonymous):

Im so sorry, i havent done this in so long, im afraid if i even tried to help, i would only make i worse! :/

OpenStudy (ness9630):

Too many questions, no one is going to want to help with 6 questions at once, even if they are simple.

OpenStudy (ness9630):

You would of been better off if you listened and followed along with shamil

OpenStudy (anonymous):

who said you had to answer all of them?

OpenStudy (anonymous):

Sorry, way too many questions.

OpenStudy (anonymous):

You don't have to answer all of them!

OpenStudy (ranga):

1) 4r^2 - 4r + 1 = 4r^2 -2r - 2r + 1 = 2r(2r - 1) - 1(2r - 1) = (2r - 1)(2r - 1) or (2r - 1)^2

OpenStudy (nincompoop):

I would read up on this one http://www.mathsisfun.com/algebra/binomial-theorem.html